Systems Staff Meeting Notes 4/6/98

Systems Staff Meeting Notes
April 6, 1998

Present: Brian Burton, Bill Covey, Debra Harris, Sharon Henry, Bobbie Parker, Suzy Shaw

Review Minutes of March 30, 1998
  1. Software #3 should be LIS tracking, not OPS.
  1. Bobbie, Debra & the students have no NT2 visible in their Network Neighborhood. Apparently have access to it, itís just not showing in the browser.
  2. Outlook will not install MSD file on Winstonís machine. Bill & Suzy had no problems - may be that it is outside the default directory.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. "Centurion Guard" ordered for work station protection. It is essentially a re-director, appearing as if it writes to the C drive, but re-directing them to the D drive. Upon reboot, it wipes out the D drive. It has a key plate on the back - a key is necessary to modify it. No restriction on files or applications. May consider using it with the staff as well - it would get rid of bad configurations and provide all storage on the network server. "Centurion Guard" runs about $80 each - free upgrade for NT. It fits on the floppy drive cable, basically, it is an internal boggle to prevent the floppy from being a boot drive. It works with any kind of hard drive & uses 9K above the 640 line. Virtual memory is sometimes slowed down. Will eliminate cache being held on a particular machine. It has been available for 5 years - FSU has about 300 of them in use. We will receive about 6 of them in for a trial run. It would be hard to beat this, unless you could get into the case. BIOS should be password protected.
  2. 57 Cows have been ordered. 20 with 17" monitors for the public. 20 with 15" monitors for the public. 17 with 17" monitors for staff. They are all 32M machines with 2 Gig HD.
  1. Routing problems have been occurring on campus, possibly due to CORE tests. Meltdowns in PowerHubs may also be the culprit.
  2. The rest of the switch purchase was supposedly shipped overnight last Friday. Has not yet arrived.
  3. Systems wiring - mix up with PPD - corresponding the phone installation. On track now, work is in progress.
  4. Preservation - info pack released to potential bidders.
  5. MSL - shelving is down, wall parts have been received.
Software Projects Status
  1. Mellon - still not finished.
  2. Personnel - Faculty & staff side is still in progress, may be ready in 6 weeks.
  3. Paula - adding the fire alarm printout button.
  4. Pending Special Collections project: migrating their patron tracking system off of the Mac. Brian will begin working on this one.
  5. Baldwin Collection - ready for 10,000 records to be uploaded to NOTIS. Winston will manipulate the records for uploading to FCLA.
  6. Microsoft Support Contract to be renewed.
  1. Sharon out today.
  2. Debra & Destry out April 10.
  3. Destry out April 13.
  4. Winston & Rich out April 13 - 20.
  5. Brian out April 16 - 17.
  6. Sharon out April 24 - 27.
  7. Memorial Day observed May 25.
  8. Bobbie out June 8 - 19.
  9. Independence Day observed July 3.
  10. Debra out July 7 - 10.
  11. Bobbie out July 13 - 17.
  12. Brian out July 16 & 17.

Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris